# Axion Dark Matter Archaeology with Primordial Gravitational Waves

### v1

@@ 3 Primordial Gravitational Waves
$\Omega_{\text{GW}}(k)=\frac{1}{12}\left(\frac{k}{a_{0}H_{0}}\right)^{2}T_{T}^{2}(k)\;P_{T}^{\text{prim}}(k),$ (3.1)
@@
with $a_{0}(H_{0})$ being the Friedmann-Lemaitre-Robertson-Walker (FLRW) metric scale factor (Hubble parameter) at the present time [63] [^1] Throughout our study $a_{0}=1$ and $H_{0}=67$ km/s/Mpc.. $P_{T}^{\text{prim}}$ represents the primordial power spectrum from inflation which can be expressed as [108, 109, 110, 111, 112, 113, 114]
$P_{T}^{\text{prim.}}(k)=A_{T}(k)\left(\frac{k}{k_{\ast}}\right)^{n_{T}}.$ (3.2)
@@
$A_{T}$ represents the amplitude of the tensor power spectra and determines the amplitude of the GW spectrum. It is evaluated at a particular pivot scale $k_{\ast}=0.05$ Mpc-1, where the Planck collaboration reports its observational constraints [92]. Additionally, $n_{T}$ is the tensor spectral index, which governs the tilts of the GW spectrum. A positive $n_{T}$ corresponds to a blue-tilted spectrum, while a negative $n_{T}$ leads to red-tilted spectrum. Although standard single-field slow-roll inflation generally predicts a red-tilted $n_{T}$ [115], several alternative scenarios can produce a blue-tilted $n_{T}$ spectrum, for instances, string cosmology [116, 117], modified gravity [118], particle production during inflation [119, 120], G-inflation [121], Higgsed chromo-natural inflation [122], natural inflation on steep potentials [123], super-inflation modes [124], Einstein-Gauss-Bonnet inflation [125], spectator scalar field during the inflation [126], Axion gauge field inflation [127, 128], Higgsed gauge-flation [129]. Furthermore, recent NANOGrav results have been interpreted as a blue-tilted spectrum, showing a good fit [130, 131].
@@
Blue-tilted spectra can overproduce gravitons, that will contribute to the additional effective number of relativistic degrees of freedom, $\Delta N_{\rm eff}$ [132]
$\int^{\infty}_{f_{\rm min}}\frac{df}{f}\Omega_{\rm GW}(f)h^{2}\;\leq 5.6\times 10^{-6}\;\Delta N_{\rm eff},$ (3.3)
[…]
@@
$T_{T}(k)$ in Eq. (3.1) is the transfer function describing the evolution of gravitational waves throughout different epochs, in the background of a FLRW Universe. In general, it can be expressed as [142, 143, 144, 50, 145, 146]:
$T_{T}^{2}(k)\equiv\Omega_{m}^{2}\left(\frac{g_{*}(T_{\text{in}})}{g_{*}^{0}}\right)\left(\frac{g_{*S}^{0}}{g_{*S}(T_{\text{in}})}\right)^{\frac{4}{3}}\left(\frac{3j_{1}(z_{k})}{z_{k}}\right)^{2}F(k),$ (3.4)
@@
where $g_{\star\,S}$ is the total number of entropy degrees of freedom and $\Omega_{m}(=0.31)$ is the total matter density [92]. $j_{1}(z_{k})$ is the first order spherical Bessel function, with $z_{k}\equiv k\tau_{0}$, where $\tau_{0}=2/H_{0}$ [63, 92] being the conformal time today. Now, the imprints of any non-standard evolution during the pre-BBN epoch are encapsulated within $F(k)$. In the standard scenario (i.e. only radiation domination after inflationary reheating and before BBN), $F(k)$ has the following form:
$F(k)\bigg{|}_{\text{standard}}\equiv T_{1}^{2}\left(\frac{k}{k_{\text{eq}}}\right)T_{2}^{2}\left(\frac{k}{k_{\text{RH}}}\right),$ (3.5)
@@
$k_{\text{eq}}$ $=$ $7.1\times 10^{-2}\Omega_{m}h^{2}\,{\rm Mpc}^{-1},$ (3.6)
$k_{\text{RH}}$ $=$ $1.7\times 10^{14}\left(\frac{g_{*S}(T_{\rm RH})}{g_{*S}^{0}}\right)^{\frac{1}{6}}\left(\frac{T_{\rm RH}}{10^{7}\;\text{GeV}}\right)\,{\rm Mpc}^{-1}.$ (3.7)
@@
$F(k)\bigg{|}_{\text{HQD}}\equiv T_{1}^{2}\left(\frac{k}{k_{\text{eq}}}\right)T_{2}^{2}\left(\frac{k}{k_{\text{dec}}}\right)T_{3}^{2}\left(\frac{k}{k_{\text{dec, S}}}\right)T_{2}^{2}\left(\frac{k}{k_{\text{RH, S}}}\right),$ (3.8)
@@
$k_{\text{dec}}$ $=1.7\times 10^{14}\left(\frac{g_{*S}(T_{\rm dec})}{g_{*S}^{0}}\right)^{\frac{1}{6}}\left(\frac{T_{\rm dec}}{10^{7}\;\text{GeV}}\right)\,{\rm Mpc}^{-1}.$ (3.9)
The characteristic wave-numbers $k_{\text{dec, S}}$ and $k_{\text{RH, S}}$, in Eq. (3.8) can be expressed as
$k_{\text{dec, S}}$$\equiv$ $k_{\text{dec}}\Delta_{s}^{2/3},$ (3.10)
$k_{\text{RH, S}}$$\equiv$ $k_{\text{RH}}\Delta_{s}^{-1/3},$ (3.11)
where $\Delta_{s}$ is the entropy dilution factor which quantifies the the change of comoving entropy density due to the decay of the heavy quark. It is defined as the ratio of comoving entropy density after HQD to that before its onset, i.e., $\Delta_{s}\equiv S(T_{\rm dec})/S(T_{\rm dom})$. To accurately determine this dilution factor, we numerically solved the coupled Friedmann Boltzmann equations, given in Eqs. (2.10), ensuring a precise treatment of the entropy evolution throughout the process.
The fitting functions ($T_{1}^{2}(x),T_{2}^{2}(x)$ and $T_{3}^{2}(x)$) in Eq. (3.4) can be expressed as [146]
$T_{1}^{2}(x)$$\equiv$ $1+1.57x+3.42x^{2},$ (3.12)
$T_{2}^{2}(x)$$\equiv$ $(1-0.22x^{3/2}+0.65x^{2})^{-1},$ (3.13)
$T_{3}^{2}(x)$$\equiv$ $1+0.59x+0.65x^{2}.$ (3.14)
@@
3.1 GW Detection prospects with interferometers
@@
    Terrestrial detectors: Laser Interferomenter Gravitational-wave Observatory (LIGO) [147, 148, 149, 150], Einstein Telescope (ET) [151, 152], Advanced LIGO (a-LIGO) [153, 154], LIGO-India [155, 156, 157, 158, 159], VIRGO [160, 161, 162, 163], advanced-VIRGO  [164, 165] [166], Cosmic Explorer (CE) [167]
@@
    Space based detectors: Deci-Hertz Interferometer Gravitational-wave Observatory (DECIGO) [168], Upgraded DECIGO (U-DECIGO) [169, 170, 171], Big-Bang Observer (BBO) [172, 173], Laser Interferometer Space Antenna (LISA) [174, 175], $\mu$-ARES [176].
@@
    Pulsar timing arrays: Square Kilometre Array (SKA) [177, 178, 179], European Pulsar Timing Array (EPTA) [180, 181, 182], North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [183, 184, 185, 186, 187]
@@
    Cosmic microwave background radiation: Detecting BB-modes in CMB with current and future satellites. Current observational constraints from Planck 2018 [92]. Future CMB experiments include CMB Bharat [134], CMB-S4 [135, 136], PICO [137], CMB-HD [138], COrE [139], the South Pole Telescope [140], and the Simons Observatory [141].
@@
Figure 3: Illustration of GW spectrum for different values of $m_{Q}$, $d$ and $T_{\rm RH}$. In each of the GW spectrum (dashed, dot-dashed, dotted), variation with respect to parameters are quoted. Grey hatched regions are the excluded regions by the existing constraints (Plank-18+BAO, BBN and LIGO). Left panel is for $n_{T}=0$, whereas Right panel is for $n_{T}=0.5$. Here $m_{Q}$ and $T_{\rm RH}$ are in GeV. For both of the plots $r=0.036$ is fixed which corresponds to the Hubble scale of inflation $H_{\rm inf}\sim 10^{13}$ GeV.
[…]
@@
Additionally, in Fig. 3, we show the power-law integrated (PLI) sensitivity curves [190] for various GW experiments, gray shaded regions are the existing bounds while rest are future sensitivty projections. These curves assume a GW spectrum varies with frequency as $\Omega_{\rm GW}\sim f^{b}$, with $b$ being the power-law exponent. The region enclosed by the PLI curve typically indicates the parameter space where a power-law signal can be detected with robust signal-to-noise ratio (SNR). Although this provides a useful benchmark, robust detection is possible only for PLI integrated curves. Therefore, we rely on a direct SNR computation to assess detectability for specific model parameters.
@@
The PLI sensitivity curves regarding gravitational wave observatories are shown in Fig. 3. In particular, we show ground-based ET [151, 152], space-based LISA [174, 175] interferometer GW missions as well as the next generation space-based experiments like BBO [172, 173] and DECIGO [168]. We also show PTAs, including the futuristic ones like the SKA [177, 178, 179] telescope.
@@
In both panels of Fig. 3, black solid line represents a benchmark parameter set with $m_{Q}=10^{14}$ GeV, $d=6$, $T_{\rm RH}=10^{16}$ GeV and $\Lambda=m_{\rm Pl}$. The other four curves (blue dashed, green dotted, red dot-dashed and purple short-long dashed) represent the variation of single parameters, annotated in the figure with corresponding colour, while keeping the others fixed. For instance, red dot-dashed line corresponds to $m_{Q}=10^{11}$ GeV while $d$, $T_{\rm RH}$ and $\Lambda$ remains unchanged from the black solid line. Both figures exhibit two distinct suppressions in the GW spectrum at different frequencies, where the lower one marks the periods of heavy quark decay and the higher one corresponds to inflationary reheating with $k_{\rm dec}<k_{\rm RH}$. The duration of HQD and the starting point of the suppression depend on both $T_{\rm dom}$ and $T_{\rm dec}$ where $T_{\rm dom}$ depends only on $m_{Q}$, while $T_{\rm dec}$ is affected by $m_{Q}$, $d$ and $\Lambda$. More quantitively, $f_{\rm dec}=k_{\text{dec}}/(2\pi)$, where $k_{\rm dec}$ is related to $T_{\rm dec}$ (Eq. (3.9)). At higher $k$ the GW spectrum is effected by entropy dilution until $f_{\rm dec,S}=k_{\text{dec, S}}/(2\pi)$ (Eq. (3.10)). During this period, as the Universe is dominated by matter, GW spectrum scales as $f^{-2}$. For instance, the black solid line in both figure exhibits a fall around $f=f_{\rm dec}\approx 3\times 10^{-5}\,{\rm Hz}$, which persists up to $f=f_{\rm dec,S}\approx 10$ Hz. The suppression of the GW spectrum at lower frequencies becomes more dominant as the duration of HQD increases which of course depends on $\Gamma_{Q}^{d}$. For a particular $m_{Q}$ and $d$, $\Gamma_{Q}^{d}$ decreases as $\Lambda$ increases (Eq. (2.8)) suppressing GW spectra, this can be seen in Fig. 3. By increasing $m_{Q}$ the decay rate increases as does $T_{\rm dom}$. However, the suppression becomes less for the higher values of $m_{Q}$ and $d$, as seen in Fig. 3 and understood from Eq. (2.8).
@@
Here it is crucial to mention that PQ symmetry breaking and associated axion cosmology also leads to the formation of topological defects like global cosmic strings and domain walls which radiate GWs (see Refs.[106, 191]) and may also collapse into Primordial Black Holes [192, 193]. However, they are only relevant when PQ symmetry breaking scale $f_{a}$ is of the order of $10^{14}$ GeV or higher. This represents the upper bound of the parameter space we investigate in this article (see Fig. 11). Moreover given the period of early HQD domination such GW signals will get further suppressed leading to almost negligible signatures compared to GW from inflation that we study here, see Ref. [194, 195] for details. Nevertheless a detailed analysis of such GW analysis combined with IGW may be interesting but is beyond the scope of the present work.
3.2 Estimation of signal-to-noise ratio (SNR)

### v2

@@ III Primordial Gravitational Waves
$\Omega_{\text{GW}}(k)=\frac{1}{12}\left(\frac{k}{a_{0}H_{0}}\right)^{2}T_{T}^{2}(k)\;P_{T}^{\text{prim}}(k),$ (14)
@@
with $a_{0}(H_{0})$ being the Friedmann-Lemaître-Robertson-Walker (FLRW) metric scale factor (Hubble parameter) at the present time Datta and Samanta (2022) [^2] Throughout our study $a_{0}=1$ and $H_{0}=67$ km/s/Mpc.. $P_{T}^{\text{prim}}$ represents the primordial power spectrum from inflation which can be expressed as Meerburg et al. (2015); Liddle and Lyth (1993a); Abbott et al. (2017a, 2007); Cabass et al. (2016); Allen and Romano (1999); Abadie et al. (2012)
$P_{T}^{\text{prim.}}(k)=A_{T}(k)\left(\frac{k}{k_{\ast}}\right)^{n_{T}}.$ (15)
@@
$A_{T}$ represents the amplitude of the tensor power spectra and determines the amplitude of the GW spectrum. It is evaluated at a particular pivot scale $k_{\ast}=0.05$ Mpc-1, where the Planck collaboration reports its observational constraints Aghanim et al. (2020). Additionally, $n_{T}$ is the tensor spectral index, which governs the tilts of the GW spectrum. A positive $n_{T}$ corresponds to a blue-tilted spectrum, while a negative $n_{T}$ leads to red-tilted spectrum. Although standard single-field slow-roll inflation generally predicts a red-tilted $n_{T}$ Liddle and Lyth (1993b), several alternative scenarios can produce a blue-tilted $n_{T}$ spectrum, for instances, string cosmology Brandenberger et al. (2007); Calcagni et al. (2014), modified gravity Fujita et al. (2019), particle production during inflation Cook and Sorbo (2012); Mukohyama et al. (2014), G-inflation Kobayashi et al. (2010), Higgsed chromo-natural inflation Adshead et al. (2016), natural inflation on steep potentials Anber and Sorbo (2010), super-inflation modes Baldi et al. (2005), Einstein-Gauss-Bonnet inflation Oikonomou (2021), spectator scalar field during the inflation Biagetti et al. (2013), Axion gauge field inflation Caldwell and Devulder (2018); Dimastrogiovanni et al. (2018), Higgsed gauge-flation Adshead and Sfakianakis (2017). Furthermore, recent NANOGrav results have been interpreted as a blue-tilted spectrum, showing a good fit Kuroyanagi et al. (2021); Afzal et al. (2023).
@@
Blue-tilted spectra can overproduce gravitons, that will contribute to the additional effective number of relativistic degrees of freedom, $\Delta N_{\rm eff}$ Maggiore (2000)
$\int^{\infty}_{f_{\rm min}}\frac{df}{f}\Omega_{\rm GW}(f)h^{2}\;\leq 5.6\times 10^{-6}\;\Delta N_{\rm eff},$ (16)
where $f_{\rm min}$ is the lower limit of integration, which is  $\simeq\;10^{-10}\,(\simeq\;10^{-18})$ Hz for BBN (CMB). Current observational constraints from Planck 2018 and BAO Aghanim et al. (2020), along with BBN Cyburt et al. (2016), place stringent upper bounds on $\Delta N_{\rm eff}$. Future CMB experiments, including CMB Bharat CMB , CMB-S4 Abazajian et al. (2019, 2015), PICO Hanany et al. (2019), CMB-HD Aiola et al. (2022), COrE Delabrouille et al. (2018), the South Pole Telescope Benson et al. (2014), and the Simons Observatory Ade et al. (2019), are expected to further refine these constraints with improved sensitivity. Among current observations, Planck 2018+BAO provides one of the most stringent upper bounds on $\Delta N_{\rm eff}$, limiting it to $0.28$ Cyburt et al. (2016). Looking ahead, CMB-HD is projected to significantly improve this constraint, reaching a sensitivity of $0.027$ Aiola et al. (2022). In this study, we incorporate both the current Planck 2018+BAO limit and the projected CMB-HD constraint to analyse their implications.
@@
$T_{T}(k)$ in Eq. (14) is the transfer function describing the evolution of gravitational waves throughout different epochs, in the background of a FLRW Universe. In general, it can be expressed as Turner et al. (1993); Chongchitnan and Efstathiou (2006); Nakayama et al. (2008); Nakayama and Yokoyama (2010); Kuroyanagi et al. (2011, 2015):
$T_{T}^{2}(k)\equiv\Omega_{m}^{2}\left(\frac{g_{*}(T_{\text{in}})}{g_{*}^{0}}\right)\left(\frac{g_{*S}^{0}}{g_{*S}(T_{\text{in}})}\right)^{\frac{4}{3}}\left(\frac{3j_{1}(z_{k})}{z_{k}}\right)^{2}F(k),$ (17)
@@
where $g_{\star\,S}$ is the total number of entropy degrees of freedom and $\Omega_{m}(=0.31)$ is the total matter density Aghanim et al. (2020). $j_{1}(z_{k})$ is the first order spherical Bessel function, with $z_{k}\equiv k\tau_{0}$, where $\tau_{0}=2/H_{0}$ Datta and Samanta (2022); Aghanim et al. (2020) being the conformal time today. Now, the imprints of any non-standard evolution during the pre-BBN epoch are encapsulated within $F(k)$. In the standard scenario (i.e. only radiation domination after inflationary reheating and before BBN), $F(k)$ has the following form:
$F(k)\bigg|_{\text{standard}}\equiv T_{1}^{2}\left(\frac{k}{k_{\text{eq}}}\right)T_{2}^{2}\left(\frac{k}{k_{\text{RH}}}\right),$ (18)
@@
$k_{\text{eq}}$ $=$ $7.1\times 10^{-2}\Omega_{m}h^{2}\,{\rm Mpc}^{-1},$ (19)
$k_{\text{RH}}$ $=$ $1.7\times 10^{14}\left(\frac{g_{*S}(T_{\rm RH})}{g_{*S}^{0}}\right)^{\frac{1}{6}}\left(\frac{T_{\rm RH}}{10^{7}\;\text{GeV}}\right)\,{\rm Mpc}^{-1}.$
@@
$F(k)\bigg|_{\text{HQD}}\equiv$ $T_{1}^{2}\left(\frac{k}{k_{\text{eq}}}\right)T_{2}^{2}\left(\frac{k}{k_{\text{dec}}}\right)\times$ (21)
$T_{3}^{2}\left(\frac{k}{k_{\text{dec, S}}}\right)T_{2}^{2}\left(\frac{k}{k_{\text{RH, S}}}\right),$
@@
$k_{\text{dec}}$ $=1.7\times 10^{14}\left(\frac{g_{*S}(T_{\rm dec})}{g_{*S}^{0}}\right)^{\frac{1}{6}}\left(\frac{T_{\rm dec}}{10^{7}\;\text{GeV}}\right)\,{\rm Mpc}^{-1}.$ (22)
The characteristic wave-numbers $k_{\text{dec, S}}$ and $k_{\text{RH, S}}$, in Eq. (21) can be expressed as
$k_{\text{dec, S}}$ $\equiv$ $k_{\text{dec}}\Delta_{s}^{2/3},$ (23)
$k_{\text{RH, S}}$ $\equiv$ $k_{\text{RH}}\Delta_{s}^{-1/3},$ (24)
where $\Delta_{s}$ is the entropy dilution factor which quantifies the the change of comoving entropy density due to the decay of the heavy quark. It is defined as the ratio of comoving entropy density after HQD to that before its onset, i.e., $\Delta_{s}\equiv S(T_{\rm dec})/S(T_{\rm dom})$. To accurately determine this dilution factor, we numerically solved the coupled Friedmann Boltzmann equations, given in Eqs. (11), ensuring a precise treatment of the entropy evolution throughout the process.
The fitting functions ($T_{1}^{2}(x),T_{2}^{2}(x)$ and $T_{3}^{2}(x)$) in Eq. (17) can be expressed as Kuroyanagi et al. (2015)
$T_{1}^{2}(x)$ $\equiv$ $1+1.57x+3.42x^{2},$ (25)
$T_{2}^{2}(x)$ $\equiv$ $(1-0.22x^{3/2}+0.65x^{2})^{-1},$ (26)
$T_{3}^{2}(x)$ $\equiv$ $1+0.59x+0.65x^{2}.$ (27)
@@
III.1 GW Detection prospects with interferometers
@@
    Terrestrial detectors: Laser Interferomenter Gravitational-wave Observatory (LIGO) Abbott et al. (2016a, b, 2017b, 2017c), Einstein Telescope (ET) M. Punturo (2010); Hild et al. (2011), Advanced LIGO (a-LIGO) Aasi et al. (2015); Abbott et al. (2021), LIGO-India Iyer et al. (2011); Souradeep et al. (2017); Unnikrishnan (2013); Sathyaprakash et al. (2012); Unnikrishnan (2024), VIRGO Abadie et al. (2010); Accadia and et al (2012); Abbott et al. (2017d, e), advanced-VIRGO  Acernese and et al (2019); Acernese et al. (2015) Akutsu et al. (2019), Cosmic Explorer (CE) Reitze et al. (2019)
@@
    Space based detectors: Deci-Hertz Interferometer Gravitational-wave Observatory (DECIGO) Yagi et al. (2011), Upgraded DECIGO (U-DECIGO) Seto et al. (2001); Kawamura et al. (2006); Yagi and Seto (2011), Big-Bang Observer (BBO) Corbin and Cornish (2006); Harry et al. (2006), Laser Interferometer Space Antenna (LISA) Amaro-Seoane et al. (2017); Baker et al. (2019), $\mu$-ARES Sesana et al. (2021).
@@
    Pulsar timing arrays: Square Kilometre Array (SKA) Carilli and Rawlings (2004); Janssen et al. (2015); Weltman et al. (2020), European Pulsar Timing Array (EPTA) Kramer and Champion (2013); Lentati et al. (2015); Babak et al. (2016), North American Nanohertz Observatory for Gravitational Waves (NANOGrav) McLaughlin (2013); Arzoumanian et al. (2018); Aggarwal et al. (2019); Brazier et al. (2019); Arzoumanian et al. (2020)
@@
    Cosmic microwave background radiation: Detecting BB-modes in CMB with current and future satellites. Current observational constraints from Planck 2018 Aghanim et al. (2020). Future CMB experiments include CMB Bharat CMB , CMB-S4 Abazajian et al. (2019, 2015), PICO Hanany et al. (2019), CMB-HD Aiola et al. (2022), COrE Delabrouille et al. (2018), the South Pole Telescope Benson et al. (2014), and the Simons Observatory Ade et al. (2019).
@@
Figure 3: Illustration of GW spectrum for different values of $m_{Q}$, $d$ and $T_{\rm RH}$, compared with the standard scenario (without HQD) via thin grey solid line. In each of the GW spectrum (dashed, dot-dashed, dotted), variation with respect to parameters are quoted. Grey hatched regions are the excluded regions by the existing constraints (Planck-18+BAO, BBN and LIGO). Left panel is for $n_{T}=0$, whereas Right panel is for $n_{T}=0.5$. Here $m_{Q}$ and $T_{\rm RH}$ are in GeV. For both of the plots $r=0.036$ is fixed which corresponds to the Hubble scale of inflation $H_{\rm inf}\sim 10^{13}$ GeV.
In Fig. 3, we illustrate the behaviour the GW spectrum in presence of the HQD, as a function of frequency, assuming $r=0.036$ 333r=0.036r=0.036 represents the present upper bound on $r$ from Planck-18 + BK18 data, measured at the pivot scale Ade et al. (2021, 2015).. The left panel of Fig. 3 corresponds to $n_{T}=0$, where the prospects of detecting the signal with current and future detectors remains low. However a blue-tilted ($n_{T}>0$) spectrum is able to enhance the detection prospects, as depicted in the right panel of Fig. 3.
@@
Additionally, in Fig. 3, we show the power-law integrated (PLI) sensitivity curves Thrane and Romano (2013) for various GW experiments, gray shaded regions are the existing bounds while rest are future sensitivity projections. These curves assume a GW spectrum varies with frequency as $\Omega_{\rm GW}\sim f^{b}$, with $b$ being the power-law exponent. The region enclosed by the PLI curve typically indicates the parameter space where a power-law signal can be detected with robust signal-to-noise ratio (SNR). Although this provides a useful benchmark, robust detection is possible only for PLI integrated curves. Therefore, we rely on a direct SNR computation to assess detectability for specific model parameters.
@@
The PLI sensitivity curves regarding gravitational wave observatories are shown in Fig. 3. In particular, we show ground-based ET M. Punturo (2010); Hild et al. (2011), space-based LISA Amaro-Seoane et al. (2017); Baker et al. (2019) interferometer GW missions as well as the next generation space-based experiments like BBO Corbin and Cornish (2006); Harry et al. (2006) and DECIGO Yagi et al. (2011). We also show PTAs, including the futuristic ones like the SKA Carilli and Rawlings (2004); Janssen et al. (2015); Weltman et al. (2020) telescope.
@@
In both panels of Fig. 3, we present the behaviour of GW spectra for the four distinct HQD parameters, compared with standard scenario i.e. absence of HQD by thin grey solid line. The black solid line represents a benchmark parameter set with $m_{Q}=10^{14}$ GeV, $d=6$, $T_{\rm RH}=10^{16}$ GeV and $\Lambda=m_{\rm Pl}$. The other four curves (blue dashed, green dotted, red dot-dashed and purple short-long dashed) represent the variation of single parameters, annotated in the figure with corresponding colour, while keeping the others fixed. For instance, red dot-dashed line corresponds to $m_{Q}=10^{11}$ GeV while $d$, $T_{\rm RH}$ and $\Lambda$ remains unchanged from the black solid line. In absence of HQD, GW spectra has only one suppression which marks the matter-dominated inflationary reheating phase. The presence of HQD phase, GW spectra exhibit an additional suppression at the lower frequencies, marking the period of heavy quark decay, whereas the higher one corresponds to the same, matter-dominated inflationary reheating with $k_{\rm dec}<k_{\rm RH}$. Thus, by suppressing the GW spectrum, the presence of HQD allows higher values of $n_{T}$ compared to the standard scenario (absence of HQD). The duration of HQD and the starting point of the suppression depend on both $T_{\rm dom}$ and $T_{\rm dec}$ where $T_{\rm dom}$ depends only on $m_{Q}$, while $T_{\rm dec}$ is affected by $m_{Q}$, $d$ and $\Lambda$. More quantitively, $f_{\rm dec}=k_{\text{dec}}/(2\pi)$, where $k_{\rm dec}$ is related to $T_{\rm dec}$ (Eq. (22)). At higher $k$ the GW spectrum is effected by entropy dilution until $f_{\rm dec,S}=k_{\text{dec, S}}/(2\pi)$ (Eq. (23)). During this period, as the Universe is dominated by matter, GW spectrum scales as $f^{-2}$. For instance, the black solid line in both figure exhibits a fall around $f=f_{\rm dec}\approx 3\times 10^{-5}\,{\rm Hz}$, which persists up to $f=f_{\rm dec,S}\approx 10$ Hz. The suppression of the GW spectrum at lower frequencies becomes more dominant as the duration of HQD increases which of course depends on $\Gamma_{Q}^{d}$. For a particular $m_{Q}$ and $d$, $\Gamma_{Q}^{d}$ decreases as $\Lambda$ increases (Eq. (9)) suppressing GW spectra, this can be seen in Fig. 3. By increasing $m_{Q}$ the decay rate increases as does $T_{\rm dom}$. However, the suppression becomes less for the higher values of $m_{Q}$ and $d$, as seen in Fig. 3 and understood from Eq. (9).
@@
Here it is crucial to mention that PQ symmetry breaking and associated axion cosmology also leads to the formation of topological defects like global cosmic strings and domain walls which radiate GWs (see Refs.Gorghetto et al. (2021); Fu et al. (2023)) and may also collapse into Primordial Black Holes Ferrer et al. (2019); Conaci et al. (2024). However, they are only relevant when PQ symmetry breaking scale $f_{a}$ is of the order of $10^{14}$ GeV or higher. This represents the upper bound of the parameter space what we investigate in this article (see Fig. 9). Moreover given the period of early HQD domination such GW signals will get further suppressed leading to almost negligible signatures compared to GW from inflation that we study here, see Ref. Ghoshal et al. (2023); Datta et al. (2025) for details. Nevertheless a detailed analysis of such GW analysis combined with IGW may be interesting but is beyond the scope of the present work.
III.2 Estimation of signal-to-noise ratio (SNR)